# Strategic Informed Trading and the Value of Private Information

### v1

@@ Appendix A The Equilibrium: General Case
(50) $\begin{split}{\mathcal{M}}={\mathcal{M}}({\mathcal{Y}})&:=P_{X|G}^{1/2}(1_{d}+{\mathcal{Y}}+{\mathcal{Y}}^{\prime})^{-1}P_{X|G}^{-1/2}P_{I},\\ \Lambda_{\iota}=\Lambda_{\iota}({\mathcal{Y}})&:=\frac{1}{\alpha_{I}}P_{I}^{-1}P_{X|G}^{1/2}(1_{d}+{\mathcal{Y}}^{\prime})P_{X|G}^{-1/2},\\ \mathcal{V}=\mathcal{V}({\mathcal{Y}})&:=P_{I}^{-1}\left(P_{X}\mu_{X}-\psi_{I,0}-P_{X|G}V({\mathcal{Y}})\right).\end{split}$
Motivated by (49) we define the market signal and its precision
(51) $\begin{split}H_{\iota}&:=G+\Lambda_{\iota}Z_{N}=X+Z_{I}+\Lambda_{\iota}Z_{N},\\ P_{U,\iota}&=P_{U,\iota}({\mathcal{Y}}):=\left(P_{I}^{-1}+\Lambda_{\iota}({\mathcal{Y}})P_{N}^{-1}\Lambda_{\iota}^{\prime}({\mathcal{Y}})\right)^{-1}.\end{split}$
By observing the price, the uniformed trader has time 00 information $\sigma(H_{\iota})$ and written as a function of $H_{\iota}$, the price is $p(H_{\iota})$ where
(52) $\begin{split}p_{\iota}(h_{\iota})&=M{\mathcal{M}}(h_{\iota}+\mathcal{V})+V.\end{split}$
As the uninformed is a price taker, his optimization problem is
$\begin{split}&\inf_{\psi\in\sigma(H_{\iota})}\mathbb{E}\left[e^{-\left(\psi_{U,0}^{\prime}p_{\iota}\left(H_{\iota}\right)+\psi^{\prime}\left(X-p_{\iota}\left(H_{\iota}\right)\right)\right)}\big{|}\sigma(H_{\iota})\right].\end{split}$
Similarly to Lemma A.1 we obtain
@@
Let $h_{\iota}\in\mathbb{R}$. On the set $\left\{H_{\iota}=h_{\iota}\right\}$, the uninformed agent has risk-aversion adjusted optimal demand
(53) $\widehat{\psi}_{U,\iota}(h_{\iota})=P_{U,\iota}h_{\iota}+P_{X}\mu_{X}-(P_{U,\iota}+P_{X})p_{\iota}(h_{\iota}).$
The market-clearing condition is
(54) $\begin{split}\Pi&=\alpha_{I}\left(\widehat{\psi}_{I,\iota}(G,Z_{N})-\psi_{I,0}+\frac{Z_{N}}{\alpha_{I}}\right)+\alpha_{I}\psi_{I,0}+\alpha_{U}\widehat{\psi}_{U,\iota}(H_{\iota}).\end{split}$
Taking into account (49), (50) and (53), the above will induce the equilibrium conditions on ${\mathcal{Y}}$ and $V$, as stated in the next proposition. To state it, as in (19) but for the general setup, define
(55) $p_{0}:=\mathbb{E}^{{\mathbb{Q}}_{0}}\left[X\right]=\mu_{X}-P_{X}^{-1}\widehat{\Pi}.$
As before, $p_{0}$ is the equilibrium price absent any private signals, and provided the initial endowments satisfy (2).
@@
(56) $\begin{split}0_{d}&=P_{U,\iota}(\widehat{{\mathcal{Y}}}){\mathcal{M}}({\mathcal{Y}})^{-1}+\frac{\alpha_{I}}{\alpha_{U}}1_{d}-(P_{U,\iota}(\widehat{{\mathcal{Y}}})+P_{X})M(\widehat{{\mathcal{Y}}}),\end{split}$
where $M$, ${\mathcal{M}}$ and $P_{U,\iota}$ are from (48), (50) and (51). Then, there exists a price-impact equilibrium. The market signal is $H_{\iota}$ from (51), and the equilibrium price $p_{\iota}$ is of the form (52) with $M=M(\widehat{{\mathcal{Y}}})$ from (48), and
(57) $V=V(\widehat{{\mathcal{Y}}})=p_{0}+P_{X}^{-1}P_{\iota}(\widehat{{\mathcal{Y}}})(P_{I}-P_{\iota}(\widehat{{\mathcal{Y}}}))^{-1}\left(\psi_{I,0}-\widehat{\Pi}\right);\qquad P_{\iota}(\widehat{{\mathcal{Y}}}):=\frac{\alpha_{I}P_{I}+\alpha_{U}P_{U,\iota}(\widehat{{\mathcal{Y}}})}{\alpha_{I}+\alpha_{U}}.$
Lastly, the price function $p_{\iota}$ takes the form
(58) $\begin{split}p_{\iota}(h_{\iota})&=p_{0}+M(\widehat{{\mathcal{Y}}}){\mathcal{M}}(\widehat{{\mathcal{Y}}})\left(h_{\iota}-p_{0}\right)\\ &\quad+\left(P_{X}^{-1}-M(\widehat{{\mathcal{Y}}}){\mathcal{M}}(\widehat{{\mathcal{Y}}})\left(P_{\iota}(\widehat{{\mathcal{Y}}})^{-1}+P_{X}^{-1}\right)\right)P_{\iota}(\widehat{{\mathcal{Y}}})(P_{I}-P_{\iota}(\widehat{{\mathcal{Y}}}))^{-1}\left(\psi_{I,0}-\widehat{\Pi}\right).\end{split}$
In light of Proposition A.3, our goal is to find solutions $\widehat{{\mathcal{Y}}}$ to (56). This is a matrix-valued cubic equation for ${\mathcal{Y}}$ and the primary difficulty in establishing existence of solutions $\widehat{{\mathcal{Y}}}$ arises due to the interaction between the precision matrices $P_{I},P_{N}$ and $P_{X}$. While techniques exist for solving such equations (see Benzi and Viviani [2023]), technically this would take us far beyond the intended scope of the paper, and therefore we make the following assumption, which is always valid in the case of a single asset (as in our simplified model of Section 1).
@@
Under this assumption we guess that ${\mathcal{Y}}=y1_{d}$ for a scalar $y>0$. The quantities in (48), (50) and (51) take the form
(59) $\begin{split}M&=\frac{y}{1+p_{I}}P_{X}^{-1},\quad{\mathcal{M}}=\frac{p_{I}}{1+2y}P_{X},\quad\Lambda_{\iota}=\frac{y+1}{\alpha_{I}p_{I}}P_{X}^{-1},\quad P_{U,\iota}=\left(\frac{\alpha_{I}^{2}p_{I}^{2}p_{N}}{(1+y)^{2}+\alpha_{I}^{2}p_{I}p_{N}}\right)P_{X}.\end{split}$
Note the constant in front of $P_{U,\iota}$ is exactly (13). Plugging these values and simplifying one can show that (56) holds if and only if $y$ solves (18). Therefore, using Proposition 1.5 we obtain the following generalization of Theorem 1.6.
@@
Let Assumption A.4 hold. Then a price-impact equilibrium exists. Using $\widehat{y}$ from Proposition 1.5, the market signal is $H_{\iota}$ from (51). The equilibrium price is $p_{\iota}(H_{\iota})$ for $p_{\iota}$ from (58). The insider uses the optimal policy $\widehat{\pi}_{I,\iota}=(1/\alpha_{I})\widehat{\psi}_{I,\iota}(G,Z_{N})$ from (49). The uninformed agent uses the optimal policy $\widehat{\pi}_{U,\iota}=(1/\alpha_{U})\widehat{\psi}_{U,\iota}(H)$ for $\widehat{\psi}_{U,\iota}$ from (53).
@@
To state the price taking results, assume there is a market signal $H$ revealed through the time 00 price $p=p(H)$, and both traders take $p(H)$ as given. The insider has time 0 information $\sigma(H,G)$ while the uninformed trader uses $\sigma(H)$. Using (46), the insider and uninformed trader’s optimal investment problems are as in (21), where again $(H,p(H))$ is a price-taking equilibrium if the clearing condition (4) holds for the optimal policies. As this result does not require Assumption A.4 we state it for general parameter values (it generalizes Proposition 1.7).
@@
(60) $H:=G+\frac{1}{\alpha_{I}}P_{I}^{-1}Z_{N}=X+Z_{I}+\frac{1}{\alpha_{I}}P_{I}^{-1}Z_{N}.$
@@
$P_{U}:=\left(P_{I}^{-1}+\frac{1}{\alpha_{I}^{2}}P_{I}^{-1}P_{N}^{-1}P_{I}^{-1}\right)^{-1}.$
With $p_{0}$ from (55), the equilibrium price is $p=p(H)$ for the price function
(61) $\begin{split}p(h)&:=p_{0}+\left(P_{X}+P\right)^{-1}P\left(h-p_{0}\right);\quad P:=\frac{\alpha_{I}P_{I}+\alpha_{U}P_{U}}{\alpha_{I}+\alpha_{U}}.\\ \end{split}$
The optimal policies for $I$ and $U$ are $\widehat{\psi}_{I}(G,H)$ and $\widehat{\psi}_{U}(H)$ respectively, where
(62) $\widehat{\psi}_{I}(g,h)=P_{X}\mu_{X}+P_{I}g-(P_{I}+P_{X})p(h);\quad\widehat{\psi}_{U}(h)=P_{X}\mu_{X}+P_{U}h-(P_{U}+P_{X})p(h).$
@@
Continuing, we state the equilibrium results (generalizing Proposition 1.11) where there is no private signal, recalling $\widehat{\Pi}$ and $p_{0}$ from (55).
@@
The no-signal equilibrium corresponds to $p_{I}=0$. In the price-taking case, the equilibrium price and the risk aversion adjusted optimal positions are $p_{ns}(Z_{N})$ and $\widehat{\psi}_{ns,I}=\widehat{\psi}_{ns,U}=\widehat{\psi}_{ns}(Z_{N})$ where
(63) $p_{ns}(z)=p_{0}+\lambda P_{X}^{-1}\frac{z}{\alpha_{I}};\qquad\widehat{\psi}_{ns}(z)=\widehat{\Pi}-\lambda\frac{z}{\alpha_{I}},$
where $\lambda$ is defined in (17). In the price-impact case, the equilibrium price is $p_{ns,\iota}(Z_{N})$ where
$\begin{split}p_{ns,\iota}(z)&=p_{0}+\frac{\lambda}{1-\lambda^{2}}P_{X}^{-1}\frac{z}{\alpha_{I}}+\frac{\lambda^{2}}{1-\lambda^{2}}P_{X}^{-1}\left(\psi_{I,0}-\widehat{\Pi}\right).\end{split}$
The risk aversion adjusted optimal policies for $I$ and $U$ are $\widehat{\psi}_{ns,\iota,I}(Z_{N})$ and $\widehat{\psi}_{ns,\iota,U}(Z_{N})$, where
$\begin{split}\widehat{\psi}_{ns,\iota,I}(z)&=\psi_{I,0}-\frac{\lambda}{1+\lambda}\frac{z}{\alpha_{I}}-\frac{1}{1+\lambda}\left(\psi_{I,0}-\widehat{\Pi}\right),\\ \widehat{\psi}_{ns,\iota,U}(z)&=\psi_{I,0}-\frac{\lambda}{1-\lambda^{2}}\frac{z}{\alpha_{I}}-\frac{1}{1-\lambda^{2}}\left(\psi_{I,0}-\widehat{\Pi}\right).\end{split}$
Appendix B Proofs from Section 1 and Appendix A
Let us first provide a mapping between the results of Section 1 and Section A. First, Lemmas 1.1, 1.3 follow directly from Lemmas A.1, A.2 respectively. Next, while Lemma 1.4 could be deduced from Proposition A.3, the algebra is quite complicated so we will provide a short stand alone proof. Proposition 1.5 requires its own proof. In Theorem 1.6, the statement regarding the price function $p_{\iota}$ follows from Proposition A.3, and that regarding $\widehat{\psi}_{U,\iota}$ follows from Lemma A.2. As for $\widehat{\psi}_{I,\iota}$ we will offer a stand-alone derivation. Lastly, Propositions 1.7, 1.11 follow from Propositions A.6, A.7 respectively. Therefore, we will first provide the necessary proofs from Section 1 and then move on to Appendix A.
Proof of Lemma 1.4.
(16) can be written
$\begin{split}0&=\alpha_{I}\left(\frac{p_{I}}{1+2y}(h_{\iota}+\widehat{\Pi})-\frac{1+p_{I}}{1+2y}(V(y)+\widehat{\Pi})\right)\\ &\qquad+\alpha_{U}\left(\left(p_{U,\iota}-\frac{(1+p_{U,\iota})p_{I}y}{(1+p_{I})(1+2y)}\right)(h_{\iota}+\widehat{\Pi})-\frac{(1+p_{U,\iota})(1+y)}{1+2y}(V(y)+\widehat{\Pi})\right)\\ &\qquad+\alpha_{U}\left(-1+\frac{(1+p_{U,\iota})y}{(1+p_{I})(1+2y)}-p_{U,\iota}+\frac{(1+p_{U,\iota})p_{I}y}{(1+p_{I})(1+2y)}+\frac{(1+p_{U,\iota})(1+y)}{1+2y}\right)\widehat{\Pi}.\end{split}$
Direct calculation shows the term on the third line above vanishes, so that if $V(y)=-\widehat{\Pi}$ the clearing condition will hold provided
(64) $\begin{split}0&=\alpha_{I}\frac{p_{I}}{1+2y}+\alpha_{U}\left(p_{U,\iota}-\frac{(1+p_{U,\iota})p_{I}y}{(1+p_{I})(1+2y)}\right),\\ &=\frac{\alpha_{I}p_{I}}{(1+2y)((1+y)^{2}+\alpha_{I}^{2}p_{I}p_{N}}\left((1+y)^{2}\left(1-\frac{\alpha_{U}y}{\alpha_{I}(1+p_{I})}\right)+\alpha_{I}^{2}p_{I}p_{N}\left(\frac{\alpha_{U}}{\alpha_{I}}y+\frac{\alpha_{U}+\alpha_{I}}{\alpha_{I}}\right)\right).\end{split}$

### v2

@@ Appendix A The Equilibrium: General Case
(51) $\begin{split}{\mathcal{M}}={\mathcal{M}}({\mathcal{Y}})&:=P_{X|G}^{1/2}(1_{d}+{\mathcal{Y}}+{\mathcal{Y}}^{\prime})^{-1}P_{X|G}^{-1/2}P_{I},\\ \Lambda_{\iota}=\Lambda_{\iota}({\mathcal{Y}})&:=\frac{1}{\alpha_{I}}P_{I}^{-1}P_{X|G}^{1/2}(1_{d}+{\mathcal{Y}}^{\prime})P_{X|G}^{-1/2}.\end{split}$
Motivated by (50) and the clearing condition (44) we define
(52) $\begin{split}H_{0,\iota}&:=G_{0}+\Lambda_{\iota}Z_{N}=X_{0}+Z_{I}+\Lambda_{\iota}Z_{N},\\ P_{U,\iota}&=P_{U,\iota}({\mathcal{Y}}):=\left(P_{I}^{-1}+\Lambda_{\iota}({\mathcal{Y}})P_{N}^{-1}\Lambda_{\iota}^{\prime}({\mathcal{Y}})\right)^{-1}.\end{split}$
Next, by observing the price, the uninformed trader has time 00 information $\sigma(H_{0,\iota})$ and from (48), (50) the de-meaned price is $p_{0,\iota}=p_{0,\iota}(H_{0,\iota})=p_{0,\iota}(H_{0,\iota})({\mathcal{Y}})$ where
(53) $\begin{split}p_{0,\iota}(h_{0,\iota})({\mathcal{Y}})&=M_{p,\iota}({\mathcal{Y}}){\mathcal{M}}({\mathcal{Y}})\left(h_{0,\iota}-P_{I}^{-1}P_{X|G}V_{0,p,\iota}({\mathcal{Y}})\right)+V_{0,p,\iota}({\mathcal{Y}}),\\ &=M_{p,\iota}({\mathcal{Y}}){\mathcal{M}}({\mathcal{Y}})h_{0,\iota}+P_{X|G}^{-1/2}(1_{d}+{\mathcal{Y}}^{\prime})(1_{d}+{\mathcal{Y}}+{\mathcal{Y}}^{\prime})^{-1}P_{X|G}^{1/2}V_{0,p,\iota}({\mathcal{Y}}).\end{split}$
With this signal and pricing function, the uninformed trader solves the problem
(54) $\begin{split}\inf_{\theta\in\sigma(H_{0,\iota})}\left(\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\theta^{\prime}\left(X_{0}-p_{0,\iota}\right)}\big{|}\sigma(H_{0,\iota})\right];\quad p_{0,\iota}=p_{0,\iota}(H_{0,\iota})\right),\end{split}$
and we have
@@
Let $h_{\iota}\in\mathbb{R}$. On $\left\{H_{0,\iota}=h_{0,\iota}\right\}$, the unique optimizer of (54) is
(55) $\widehat{\theta}_{U,\iota}(h_{0,\iota})=P_{U,\iota}h_{0,\iota}-(P_{U,\iota}+P_{X})p_{0,\iota}(h_{0,\iota}).$
In view of (52), the market-clearing condition (44) can be rewritten
(56) $\begin{split}0&=\alpha_{I}\left(\widehat{\theta}_{I,\iota}(G_{0},Z_{N})+\frac{Z_{N}}{\alpha_{I}}\right)+\alpha_{U}\widehat{\theta}_{U,\iota}(H_{0,\iota}),\\ &=\alpha_{I}{\mathcal{M}}({\mathcal{Y}})\left(H_{0,\iota}-P_{I}^{-1}P_{X|G}V_{0,p,\iota}({\mathcal{Y}})\right)\\ &\qquad\qquad+\alpha_{U}\left(P_{U,\iota}({\mathcal{Y}})H_{0,\iota}-(P_{U,\iota}({\mathcal{Y}})+P_{X})p_{\iota}(H_{0,\iota})({\mathcal{Y}})\right).\end{split}$
Using (53) it is clear that $V_{0,p,\iota}({\mathcal{Y}})=0$. To eliminate the $H_{0,\iota}$ terms we need
(57) $0_{d}=\alpha_{I}{\mathcal{M}}({\mathcal{Y}})+\alpha_{U}P_{U,\iota}({\mathcal{Y}})-\alpha_{U}(P_{U,\iota}({\mathcal{Y}})+P_{X})M_{p,\iota}({\mathcal{Y}}){\mathcal{M}}({\mathcal{Y}}).$
Re-writing this equation slightly, and translating back to the un-demeaned values, we obtain
@@
(58) $\begin{split}0_{d}&=\frac{\alpha_{I}}{\alpha_{U}}1_{d}+P_{U,\iota}(\widehat{{\mathcal{Y}}}){\mathcal{M}}({\mathcal{Y}})^{-1}-(P_{U,\iota}(\widehat{{\mathcal{Y}}})+P_{X})M_{p,\iota}(\widehat{{\mathcal{Y}}}),\end{split}$
where $M_{p,\iota}$, ${\mathcal{M}}$, $P_{U,\iota}$ are from (49), (51), (52). Then, there is a price-impact equilibrium. The market signal is
(59) $H_{\iota}=H_{0,\iota}+p_{{\mathbb{Q}}_{0}}=G+\Lambda_{\iota}(\widehat{{\mathcal{Y}}})Z_{N},$
where $\Lambda_{\iota}$ is from (51). The equilibrium price is
(60) $\begin{split}p_{\iota}(H_{\iota})&=p_{{\mathbb{Q}}_{0}}+M_{p,\iota}(\widehat{{\mathcal{Y}}}){\mathcal{M}}(\widehat{{\mathcal{Y}}})\left(H_{\iota}-p_{{\mathbb{Q}}_{0}}\right).\end{split}$
The risk-aversion adjusted optimal policies are
(61) $\begin{split}\widehat{\psi}_{I,\iota}&=\widehat{\Pi}+P_{X|G}^{1/2}(1+\widehat{{\mathcal{Y}}})^{-1}P_{X|G}^{-1/2}P_{X}p_{{\mathbb{Q}}_{0}}+\frac{1}{\alpha_{I}}\Lambda_{\iota}(\widehat{{\mathcal{Y}}})^{-1}G\\ &\qquad-P_{X|G}^{1/2}(1+\widehat{{\mathcal{Y}}}^{\prime})^{-1}P_{X|G}^{1/2}p_{\iota}(H_{\iota}),\\ \widehat{\psi}_{U,\iota}&=\widehat{\Pi}+P_{X}p_{{\mathbb{Q}}_{0}}+P_{U,\iota}H_{\iota}-(P_{U,\iota}+P_{X})p_{\iota}(H_{\iota}).\end{split}$
In light of Proposition A.3, our goal is to find solutions $\widehat{{\mathcal{Y}}}$ to (58). This is a matrix-valued cubic equation for ${\mathcal{Y}}$ and the primary difficulty in establishing existence of solutions $\widehat{{\mathcal{Y}}}$ arises due to the interaction between the precision matrices $P_{I},P_{N}$ and $P_{X}$. While techniques exist for solving such equations (see Benzi and Viviani [2023]), technically this would take us far beyond our intended scope, and therefore we make the following assumption, which is always valid in the case of a single asset.
@@
Under this assumption we guess ${\mathcal{Y}}=y1_{d}$ for a scalar $y>0$. The quantities in (49), (51) and (52) take the form
(62) $\begin{split}M_{p,\iota}&=\frac{y}{1+p_{I}}P_{X}^{-1},\hskip 5.0pt{\mathcal{M}}=\frac{p_{I}}{1+2y}P_{X},\hskip 5.0pt\Lambda_{\iota}=\frac{y+1}{\alpha_{I}p_{I}}P_{X}^{-1},\hskip 5.0ptP_{U,\iota}=p_{U,\iota}P_{X},\end{split}$
where using the notation of (7) we have
(63) $p_{U,\iota}:=\left(\frac{\kappa p_{I}^{2}}{(1+y)^{2}+\kappa p_{I}}\right).$
By plugging in these values and simplifying one can show that (58) holds if and only if $y$ solves (6). Indeed
$\begin{split}&P_{U,\iota}(\widehat{{\mathcal{Y}}}){\mathcal{M}}({\mathcal{Y}})^{-1}+\frac{\alpha_{I}}{\alpha_{U}}1_{d}-(P_{U,\iota}(\widehat{{\mathcal{Y}}})+P_{X})M_{p,\iota}(\widehat{{\mathcal{Y}}})\\ &\qquad=\left(\frac{\kappa p_{I}(1+2y)}{(1+y)^{2}+\kappa p_{I}}+\frac{\lambda}{1-\lambda}-\left(1+\frac{\kappa p_{I}^{2}}{(1+y)^{2}+\kappa p_{I}}\right)\frac{y}{1+p_{I}}\right)1_{d}\end{split}$
The quantity within the parentheses simplifies to
$\begin{split}\frac{\lambda}{(1-\lambda)((1+y)^{2}+\kappa p_{I})}\left((1+y)^{2}\left(1-\frac{1-\lambda}{\lambda(1+p_{I})}y\right)+\frac{\kappa p_{I}}{\lambda}\left((1-\lambda)y+1\right)\right),\end{split}$
and the term within the parentheses on the right is exactly the right side of (6). Therefore, using Proposition 1.1 we obtain the following generalization of Theorem 1.2.
@@
Let Assumption A.4 hold. Then a price-impact equilibrium exists. Using $\widehat{y}$ from Proposition 1.1, the market signal is $H_{\iota}$ from (59). The equilibrium price is $p_{\iota}(H_{\iota})$ from (60). The optimal policies for the insider and uninformed trader are from (61).
@@
For the price-taking results, assume there is a market signal $H$ revealed by the time 00 price $p=p(H)$, and both traders take $p(H)$ as given. The insider has time 0 information $\sigma(H,G)$ while the uninformed trader uses $\sigma(H)$. We make the same translations as in the price-impact case (with the obvious notation) so that the insider and uninformed trader’s optimal investment problems are
(64) $\begin{split}(I)&\qquad\inf_{\theta\in\sigma(G_{0},H_{0})}\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\theta^{\prime}(X_{0}-p_{0})}\big{|}\sigma(G_{0},H_{0})\right],\\ (U)&\qquad\inf_{\theta\in\sigma(H_{0})}\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\theta^{\prime}(X_{0}-p_{0})}\big{|}\sigma(H_{0})\right].\end{split}$
The clearing condition is again (44). Provided we find a market-clearing signal $H_{0}$ and price $p_{0}$ for the above optimization problems, the equilibrium signal and price will be $H=H_{0}+p_{{\mathbb{Q}}_{0}}$ and $p=p_{0}+p_{{\mathbb{Q}}_{0}}$ respectively, and the optimal risk aversion adjusted policies will be $\widehat{\psi}_{i}=\widehat{\theta}_{i}+\widehat{\Pi},i\in\left\{I,U\right\}$. As this result does not require Assumption A.4 we state it for general parameter values in the following.
@@
(65) $H=H_{0}+p_{{\mathbb{Q}}_{0}}=G_{0}+\frac{1}{\alpha_{I}}P_{I}^{-1}Z_{N}+p_{{\mathbb{Q}}_{0}}=G+\frac{1}{\alpha_{I}}P_{I}^{-1}Z_{N}.$
@@
(66) $P_{U}:=\left(P_{I}^{-1}+\frac{1}{\alpha_{I}^{2}}P_{I}^{-1}P_{N}^{-1}P_{I}^{-1}\right)^{-1}.$
The equilibrium price is
(67) $\begin{split}p(H)&=p_{{\mathbb{Q}}_{0}}+\left(\alpha_{I}(P_{I}+P_{X})+\alpha_{U}(P_{U}+P_{X})\right)^{-1}(\alpha_{I}P_{I}+\alpha_{U}P_{U})\left(H-p_{{\mathbb{Q}}_{0}}\right).\end{split}$
The optimal policies for $I$ and $U$ are
(68) $\begin{split}\widehat{\psi}_{I}&=\widehat{\Pi}+P_{X}p_{{\mathbb{Q}}_{0}}+P_{I}G-(P_{I}+P_{X})p(H),\\ \widehat{\psi}_{U}&=\widehat{\Pi}+P_{X}p_{{\mathbb{Q}}_{0}}+P_{U}H-(P_{U}+P_{X})p(H).\end{split}$
@@
We next state the equilibrium results, generalizing Proposition 1.4) for no private signal. These results are valid under Assumption A.4. Here, with the notation of (7) $P_{U}$ from (66) reduces to (compare with (63))
(69) $P_{U}=p_{U}P_{X};\qquad p_{U}:=\left(\frac{\kappa p_{I}^{2}}{1+\kappa p_{I}}\right)P_{X},$
which shows that $p_{U}\to 0$ as $p_{I}\to 0$ on the order of $p_{I}^{2}$.
@@
Under Assumption A.4, the no-signal equilibrium corresponds to $p_{I}=0$. In the price-taking case, the equilibrium price and the risk aversion adjusted optimal positions are
(70) $p_{ns}=p_{{\mathbb{Q}}_{0}}+\lambda P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}};\qquad\widehat{\psi}_{ns,I}=\widehat{\psi}_{ns,U}=\widehat{\Pi}-\lambda\frac{Z_{N}}{\alpha_{I}},$
where $\lambda$ is defined in (7). In the price-impact case, the equilibrium price is
$\begin{split}p_{ns,\iota}&=p_{{\mathbb{Q}}_{0}}+\frac{\lambda}{1-\lambda^{2}}P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}}.\end{split}$
The risk aversion adjusted optimal policies for $I$ and $U$ are
$\begin{split}\widehat{\psi}_{ns,\iota,I}&=\widehat{\Pi}-\frac{\lambda}{1+\lambda}\frac{Z_{N}}{\alpha_{I}};\qquad\widehat{\psi}_{ns,\iota,U}=\widehat{\Pi}-\frac{\lambda}{1-\lambda^{2}}\frac{Z_{N}}{\alpha_{I}}.\end{split}$
Appendix B Proofs
B.1. Proofs from Section 1
In this section, we first prove the cubic equation result Proposition 1.1. We then translate the general equilibrium results of Theorem A.5, Proposition A.6 and Proposition A.7 into the forms given in Theorem 1.2, Proposition 1.3 and Proposition 1.4 respectively.